40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 104     Entry time: Tue Aug 23 09:44:11 2016
Author: Gabriele 
Type: General 
Category: General 
Subject: Measurements on MO-02 

Two good ring-downs measurements were performed on MO-02. The first one was already reported in a previous elog entry. I performed another measurement, and refined the mode identification. I think I had misidentified some modes in my previous analysis. The following plot shows the difference between the modes as predicted by COMSOL and as measured. A clean quadratic trend is visible and fitted:

Here's the spectrum with all the modes:

And the updated Q measurement plot:

A second ring down was measured on Monday morning . Here are the relevant plots:

This is the same disk as before, but almost all Q values are systematically higher. Here's a direct comparison:

I'm not sure what changed between the two measurements, except for a re-alignment of the QPD. The disk might have moved a bit...

Attachment 4: spectrum_with_modes_12_01pm.png  161 kB  | Hide | Hide all
spectrum_with_modes_12_01pm.png
ELOG V3.1.3-